# Entity-based Reinforcement Learning for Autonomous Cyber Defence

### v2

@@ 3.1 Entity Gym
The Environment class in Entity Gym defines a standard interface for entity-based RL. Following a similar structure to Gymnasium (Towers et al.,, 2024), this interface includes methods for defining the observation and action spaces, resetting the environment, processing actions, and generating observations. To create a custom entity-based environment, one inherits from the Environment class and implements these core abstract methods. This process involves defining all relevant entity types and their features, as well as the action spaces associated with specific entity types. As we do in this paper, if replicating an existing environment such as Yawning Titan which has a Gym or Gymnasium interface, it’s important to ensure that the observations and actions in the Entity Gym interface correspond to those in the original version and that functionality is preserved. In other words, it should be the case that the interface and representation of the environment is different, but the information provided to the agent and the fundamental dynamics of the environment are the same.
@@
Any entity-based environment is initialised with lists of possible entities, all with a specified type. Each entity type is simply defined by a list of named features that are associated with that entity type. Observations
Each observation can contain global features, entity-specific features, action masks and rewards for the current timestep. The entity features are stored in a dictionary mapping entity types to lists of feature arrays, where each element in the list are the features for a particular entity. These lists are permitted to be any length at any given timestep, allowing for variable numbers of entities in the environment. Action masks, defined for each action type, specify which entities can perform actions and what actions are available at a particular timestep. These masks directly interact with the action space, constraining the set of valid actions for each entity. We don’t make use of action masks for our environment, although one might imagine them being useful for a real cyber environment, where access to particular nodes may be prevented.
@@
  • GlobalCategoricalActionSpace’: For global actions that aren’t associated with specific entities.
@@ 3.2 RogueNet
RogueNet takes ‘entity-based’ observations as input, represented as a set of features for each entity. The input is structured as a mapping from entity types to ragged batch objects, allowing for efficient training with batching whilst maintaining support for variable numbers of entities across environment instances. These batches are constructed from the lists of entity features that constitute the observations provided to the agent at each timestep. Masks are used to prevent information leakage between timesteps and environment instances. Global features, such as network connectivity information, could be incorporated and are appended to each entity’s feature set, maintaining a consistent input structure.
@@
After entities are embedded, these embeddings are passed through several layers of Transformer blocks, in our case 2. Each block comprises self-attention and feedforward sublayers with residual connections and layer normalisation.
@@
RogueNet can be trained using policy gradient methods, and we use the version of PPO implemented in the companion Entity Neural Network Trainer package (Winter et al., 2023c, ), which is itself adapted from CleanRL (Huang et al., 2022b, ). RogueNet’s use of ragged tensors and the attention mechanism enables it to process varying numbers of entities without altering its components or structure. This property allows the architecture to be applied to environments with dynamic entity populations or across multiple environments with different numbers of entities, such as network environments with different numbers of nodes. It is straightforward to deploy a trained policy on any Entity Gym environment containing the same types of entities and action spaces as the environment used for training the policy. For our use-case, this is network environments containing variable numbers of nodes.
@@ 4 Yawning Titan
Yawning Titan is an abstract, graph-based environment designed for the development of RL agents for autonomous cyber defence (Andrew et al.,, 2022). It provides an OpenAI Gym (Brockman et al.,, 2016) interface for a network-based cyber defence game through which a defending agent (blue agent) can interact with a network environment containing an attacker (red agent). Yawning Titan allows for flexibility in the configuration of the network and the dynamics of the game, including types of nodes, network topology, actions and observations available, termination, rewards and red agent behaviour. Several works have used and adapted Yawning Titan to train and evaluate autonomous blue agents (Acuto and Maskell,, 2023; Andrew et al.,, 2022; Collyer et al.,, 2022) with various different environment configurations. In the following, we describe the main components of the environment configuration we have used for the experiments in this paper, following previous work closely (Collyer et al.,, 2022) but with some differences to allow for clearer comparisons. Yawning Titan includes more features and options for configuration than discussed here. More information can be found in the documentation and repository, as well as the original paper (Andrew et al.,, 2022).
@@ 4.1 Nodes and Network
The nodes in the environment are simple objects with two main attributes, Node Vulnerability and Compromised Status. Node Vulnerability is a value in the range $[0,1]$ that dictates how easily a node can be compromised by an attacking red agent, with a higher value meaning the node is more vulnerable. Each node might also have an additional special attribute:
@@
For our experiments we used a single randomly selected Entry Node, and we did not use the High Value Node and Deceptive Node functionality. We mention them here as these are examples of possible varieties of entity type that could be included in more complex environments to make full use of the entity-based framework. Adding the deceptive node functionality would also mean that the number of nodes in the environment could vary during episodes, not just between episodes as we have tested here. This is something that a GNN or Transformer-based policy such as RogueNet should also be robust to in principle.
The network structures underlying each instance of a Yawning Titan environment are constructed using the NetworkX package (Hagberg et al.,, 2008). Following Collyer et al., (2022), we generate these networks randomly using the Erdős-Renyi model with edge probability equal to $0.1$, with post-processing to make sure the graph is connected. Figure 1 shows examples of networks generated in this way for 10, 20 and 40 nodes, which are the network sizes we used for our experiments.
@@ 4.3 Red Agent
In cyber defence simulators, the behaviour of the attacking red agent is typically generated in a rules-based fashion, either with some randomisation or entirely deterministically (Standen et al., 2022b, ; Andrew et al.,, 2022; Standen et al., 2021a, ; Standen et al., 2022a, ). The red agent is designed to spread through a given network and compromise nodes, while the blue agent is tasked with preventing this from happening.
@@
  • Zero Day attacks allow the red agent to compromise a targeted node regardless of its own skill level or the vulnerability of the node. This gives the red agent the chance to make progress even when facing a strong defensive blue agent. However, so the basic attack is not redundant, this only becomes available once every three environment steps.
@@ 5 Entity-based adaptation
In order to apply entity-based RL to Yawning Titan, we first adapted its underlying simulator to the Entity Gym interface. In this section we will provide a description of how this can be achieved.
@@ 5.1 Entity Types
In the Yawning Titan environment, it is possible to make the distinction between different types of nodes. Whilst all nodes behave in fundamentally the same way, there may be some nodes that have a special designation. These are the entry, high-value and deceptive nodes. For our experiments we used a single randomly-chosen entry node, with all others being generic. For this reason, we have chosen to use a single generic node Entity Type in the construction of the entity-based environment. In the RogueNet architecture (Winter et al.,, 2021), this means that a single shared entity encoder is used for all node embeddings during training.
@@ 5.3 Action Space
In Yawning Titan, the blue agent may only take action to intervene on one node per time-step. The Entity-Gym framework offers three possible types of action, as outlined in Section 3.1. Using the entity-level CategoricalActionSpace means that a candidate action is chosen for every entity of a particular type at every environment step. In the case of games containing a large number of simultaneously acting entities like Starcraft, this is necessary to allow the agent to act on any subset of them simultaneously, in a tractable way. In our setting this would mean that the blue agent could direct any subset of nodes simultaneously to either reset themselves or reduce their vulnerability (with any combination of these actions for the selected nodes). This option may be desirable in real network security applications, but it disrupts the balance of the Yawning Titan game, making it significantly easier for the blue agent. It is likely this would be a similar issue in other autonomous cyber defence environments. This means it is necessary to create a two-stage, structured action space using a combination of a SelectEntityActionSpace and either a CategoricalActionSpace or a GlobalCategoricalActionSpace.
@@
For selecting nodes to act on, we use a SelectEntityActionSpace, which requires defined actor and actee entity types. In order to use this action type, we create a dummy ‘Defender’ entity with fixed features to serve as an actor, which is able to act on all nodes in the network.
@@
For the node-specific actions of Reduce Vulnerability and Restore Node, these can be implemented using either the ‘global’ categorical action space or the ‘entity-level’ categorical action space. For a GlobalCategoricalActionSpace, the RogueNet architecture creates a dummy global entity with random features, with information from the nodes of the network shared with its embedding vector through the attention mechanism in the Transformer layers. This single, compressed embedding is then provided as the input for the action head. Combined with the SelectEntityActionSpace, this means the agent chooses an ‘action type’ and then a node on which to execute it. Under the entity-level CategoricalActionSpace, all embeddings of the nodes are passed into the action head simultaneously in a batch, with a separate action selected for every node independently, based on each node’s respective embedding. Conceptually, combining this with a SelectEntityActionSpace would be like choosing a node to execute the action that it has selected, rather than choosing the action centrally. In our case, we chose to use the global action space, as this should offer marginally more efficient inference due to fewer actions being calculated. Incidentally, the underlying mechanism is very similar to the approach outlined in Janisch et al., 2023b , where environments are represented as a graph, and an extra ‘global node’ is instantiated to gather global information over a series of GNN message-passing steps. This global node is then used to choose the ‘action identifier’, or action type, before proceeding in an autoregressive fashion.
@@ 6 Experiments
In this section, we outline the methodology and present results from the experiments we conducted to compare a baseline MLP-based Proximal Policy Optimization (PPO) agent (Schulman et al.,, 2017) with an entity-based RogueNet (Winter et al., 2023a, ) policy, as well as experiments exploring the zero-shot generalisation abilities of entity-based agents in environments with network topologies not seen during training. We drew the baseline PPO agent implementation from Stable Baselines 3 (Raffin et al.,, 2021), which is a library that collects stable and performant implementations of popular deep RL algorithms, and is generally reported to have the most consistent performance among open-source PPO implementations (Huang et al.,, 2024; Huang et al., 2022a, ). We trained the RogueNet agents using the companion Entity Neural Network Trainer package (Winter et al., 2023c, ) which uses a simple implementation of PPO derived from the CleanRL library(Huang et al., 2022b, ).
Figure 2: Train-time episodic rewards evaluated on the 10, 20, and 40 node networks, respectively. The four agents compared in each evaluation are: baseline PPO agent on the static (sb3_static_[nodes]) and random (sb3_random_[nodes]) network environments, and the entity neural network agent on the same static (sb3_random_[nodes]) and random (‘Entity_random_[nodes]’) environments. Rewards are averaged as the mean over three different random seeds, and shaded error bands are constructed between the maximum and minimum of the three runs.
@@ 6.1 Training
As suggested in Mern et al., (2020, 2019), one advantage of the entity-based approach and using a Transformer policy is ‘object exchangeability’. That is, the robustness of Transformers and GNNs to different permutations of objects in the observation vector or action space without the effectiveness of the policy being impacted. To demonstrate this property, we consider two different training settings, which we refer to as Static and Random. In the Static setting, a single network is randomly initialised at the start of training, and remains the same for every episode; the only thing that varies is the identity of the entry node. In the Random setting, an entirely new network is initialised for every episode. The aim of this setup is to evaluate policy robustness across inhomogeneous networks in the training phase, and is similar to the method used by Hong et al., (2022) in ‘Multitask Inhomogeneous Reinforcement Learning’ for robotic control.
@@
To produce the training graphs we used the default PPO hyperparameters as specified in the libraries for both agents. It’s worth noting that the default learning rate for the optimiser for the entity-based policy is 0.005, compared to 0.0003 in the Stable Baselines 3 implementation. Different choices of learning rate are likely to affect the gap between asymptotic reward of the two varieties of policy in the Static training regime.
@@
In total, we trained policies from 3 random seeds for each combination of policy type and environment setting. Considering all the combinations of agents, network sizes and ‘Random’ or ‘Static’ environments, we have trained 36 policies:
  • sb3_static_$n$_node_network ×$\times 3$: Three different MLP policies trained using the Stable Baselines 3 PPO implementation, on static networks of 10, 20, and 40 nodes respectively. Each policy was initialised and trained using a different random seed.
@@
  • sb3_random_$n$_node_network ×$\times 3$: Three different MLP policies trained using the Stable Baselines 3 PPO implementation, on $n$-node networks randomly generated with each episode reset for $n\in(10,20,40)$ (the Random) environment). Each policy was initialised and trained using a different random seed.
  • EntityYT_static_$n$_node_network ×$\times 3$: Three different Entity Neural Network RogueNet policies, trained using the entity-based version of Yawning Titan, on static networks of 10, 20, and 40 nodes respectively. Each policy was initialised and trained using a different random seed.
@@
  • EntityYT_random_$n$_node_network ×$\times 3$: Three different Entity Neural Network RogueNet policies, trained using the entity-based version of Yawning Titan, on $n$-node networks randomly generated with each episode reset for $n\in(10,20,40)$ (the Random) environment). Each policy was initialised and trained using a different random seed.
@@
Figure 2(a) shows the reward trajectories of all four types of policies trained on 10-node networks. Both entity-based agents (‘Entity_random_10’ and ‘Entity_static_10’) show similar reward trajectories that converge rapidly and plateau close to the maximum possible reward of 100. The Stable Baselines PPO agents trained on the single static network (‘sb3_static_10’) converged quickly as well, to a slightly lower asymptotic reward. In line with expectations, the baseline PPO agents trained on a series of random networks that vary across episodes (‘sb3_random_10’) struggled to learn a good policy and displayed high variance in their performance. This supports the notion that a Transformer entity-based policy is more robust to exchangeable representations, and performs no worse than when trained on a static environment, whereas a fixed MLP-based policy struggles to learn an effective policy when the object that each element of the agent’s observation vector represents changes with each episode (Mern et al.,, 2020, 2019; Zhao et al.,, 2022).
@@
Figures 2(b) and (c) show similar results on 20-node and 40-node networks respectively. Here, the difference between the asymptotic episodic reward of the policies trained in the Static regime grows smaller with a larger number of nodes. This may be, in part, due to the nature of the reward function defined as the proportion of non-compromised nodes at time $t$. Since the skill and behaviour of the red agent are unchanged, this naturally means that a lower proportion of nodes will be compromised on a given timestep, and thus rewards for the blue agent will be higher in larger environments. Nevertheless, the Stable Baselines agent trained on randomly generated networks still does comparatively poorly, and struggles to learn a policy that is invariant to different network topologies.

### v3

@@ 3.1 Entity Gym
The Environment class in Entity Gym defines a standard interface for entity-based RL. Following a similar structure to Gymnasium (Towers et al.,, 2024), this interface includes methods for defining the observation and action spaces, resetting the environment, processing actions, and generating observations. To create a custom entity-based environment, one inherits from the Environment class and implements these core abstract methods. This process involves defining all relevant entity types and their features, as well as the action spaces associated with specific entity types. As we do in this paper, if replicating an existing environment such as Yawning Titan which has a Gym or Gymnasium interface, it is important to ensure that the observations and actions in the Entity Gym interface correspond to those in the original version and that functionality is preserved. In other words, it should be the case that the interface and representation of the environment are different, but the information provided to the agent and the fundamental dynamics of the environment are the same.
@@
Any entity-based environment is initialised with lists of possible entities, each with a specified type. Each entity type is simply defined by a list of named features that are associated with that entity type.
Observations
Each observation can contain global features, entity-specific features, action masks, and rewards for the current time step. The entity features are stored in a dictionary mapping entity types to lists of feature arrays, where each element in the list is the feature vector for a particular entity. These lists are permitted to be any length at any given time step, allowing for variable numbers of entities in the environment. Action masks, defined for each action type, specify which entities can perform actions and which actions are available at a particular time step. These masks interact directly with the action space, constraining the set of valid actions for each entity. We do not make use of action masks for our environment, although one might imagine them being useful for a real cyber environment, where access to particular nodes may be prevented.
@@
  • GlobalCategoricalActionSpace’: For global actions that are not associated with specific entities.
@@ 3.2 RogueNet
RogueNet takes entity-based observations as input, represented as a set of features for each entity. The input is structured as a mapping from entity types to ragged batch objects, allowing for efficient training with batching whilst maintaining support for variable numbers of entities across environment instances. These batches are constructed from the lists of entity features that constitute the observations provided to the agent at each time step. Masks are used to prevent information leakage between time steps and environment instances. Global features, such as network connectivity information, could be incorporated and are appended to each entity’s feature set, maintaining a consistent input structure.
@@
After entities are embedded, these embeddings are passed through several layers of Transformer blocks, in our case 2. Each block comprises self-attention and feedforward sub-layers with residual connections and layer normalisation.
@@
RogueNet can be trained using policy gradient methods, and we use the version of PPO implemented in its companion Entity Neural Network Trainer package (Winter et al., 2023c, ), which is itself adapted from CleanRL (Huang et al., 2022b, ). RogueNet’s use of ragged tensors and the attention mechanism enables it to process varying numbers of entities without altering its components or structure. This property allows the architecture to be applied to environments with dynamic entity populations or across multiple environments with different numbers of entities, such as network environments with different numbers of nodes. It is straightforward to deploy a trained policy on any Entity Gym environment containing the same types of entities and action spaces as the environment used for training the policy. For our use-case, this is network environments containing variable numbers of nodes.
@@ 4 Yawning Titan
Yawning Titan is an abstract, graph-based environment designed for the development of RL agents for autonomous cyber defence (Andrew et al.,, 2022). It provides an OpenAI Gym (Brockman et al.,, 2016) interface for a network-based cyber defence game through which a defending agent (blue agent) can interact with a network environment containing an attacker (red agent). Yawning Titan allows for flexibility in the configuration of the network and the dynamics of the game, including types of nodes, network topology, actions and observations available, termination, rewards, and red agent behaviour. Several works have used and adapted Yawning Titan to train and evaluate autonomous blue agents (Acuto and Maskell,, 2023; Andrew et al.,, 2022; Collyer et al.,, 2022) with various different environment configurations. In the following, we describe the main components of the environment configuration we have used for the experiments in this paper, following previous work closely (Collyer et al.,, 2022) but with some differences to allow for clearer comparisons. Yawning Titan includes more features and options for configuration than discussed here. More information can be found in the documentation and repository, as well as the original paper (Andrew et al.,, 2022).
@@ 4.1 Nodes and Network
The nodes in the environment are simple objects with two main attributes, Node Vulnerability and Compromised Status. Node Vulnerability is a value in the range $[0,1]$ that dictates how easily a node can be compromised by an attacking red agent, with a higher value meaning that the node is more vulnerable. Each node might also have an additional special attribute:
@@
For our experiments, we used a single randomly selected Entry Node, and we did not use the High Value Node and Deceptive Node functionality. We mention them here as these are examples of possible varieties of entity type that could be included in more complex environments to make full use of the entity-based framework. Adding the deceptive node functionality would also mean that the number of nodes in the environment could vary during episodes, not just between episodes, as we have tested here. This is something a GNN or Transformer-based policy such as RogueNet should also be robust to in principle.
The network structures underlying each instance of a Yawning Titan environment are constructed using the NetworkX package (Hagberg et al.,, 2008). Following Collyer et al., (2022), we randomly generate these networks using the Erdős-Renyi model with edge probability equal to $0.1$, with post-processing to make sure the graph is connected. Figure 1 shows examples of networks generated in this way for 10, 20 and 40 nodes, which are the network sizes we used for our experiments.
@@ 4.3 Red Agent
In cyber defence simulators, the behaviour of the attacking red agent is typically generated in a rules-based fashion, either with some randomisation or entirely deterministic (Standen et al., 2022b, ; Andrew et al.,, 2022; Standen et al., 2021a, ; Standen et al., 2022a, ). The red agent is designed to spread through a given network and compromise nodes, while the blue agent is tasked with preventing this from happening.
@@
  • Zero Day attacks allow the red agent to compromise a targeted node regardless of its own skill level or the vulnerability of the node. This gives the red agent the opportunity to make progress even when facing a strong defensive blue agent. However, so the basic attack is not redundant, this only becomes available once every three environment steps.
@@ 5 Entity-based adaptation
In order to apply entity-based RL to Yawning Titan, we first adapted its underlying simulator to the Entity Gym interface. In this section, we will provide a description of how this can be achieved.
@@ 5.1 Entity Types
In the Yawning Titan environment, it is possible to make the distinction between different types of nodes. Whilst all nodes behave in fundamentally the same way, there may be some nodes that have a special designation. These are the entry, high-value, and deceptive nodes. For our experiments, we used a single randomly chosen entry node, with all others being generic. For this reason, we have chosen to use a single generic node Entity Type in the construction of the entity-based environment. In the RogueNet architecture (Winter et al.,, 2021), this means that a single shared entity encoder is used for all node embeddings during training.
@@ 5.3 Action Space
In Yawning Titan, the blue agent may only take action to intervene on one node per time step. The Entity Gym framework offers three possible types of action, as outlined in Section 3.1. Using the entity-level CategoricalActionSpace means that a candidate action is chosen for every entity of a particular type at every environment step. In the case of games containing a large number of simultaneously acting entities like Starcraft, this is necessary to allow the agent to act on any subset of them simultaneously, in a tractable way. In our setting, this would mean that the blue agent could direct any subset of nodes simultaneously to either reset themselves or reduce their vulnerability (with any combination of these actions for the selected nodes). This option may be desirable in real network security applications, but it disrupts the balance of the Yawning Titan game, making it significantly easier for the blue agent. It is likely this would be a similar issue in other autonomous cyber defence game environments. This means that it is necessary to create a structured two-stage action space using a combination of a SelectEntityActionSpace and either a CategoricalActionSpace or a GlobalCategoricalActionSpace.
@@
For selecting nodes to act on, we use a SelectEntityActionSpace, which requires defined actor and actee entity types. In order to use this action type, we create a dummy ‘Defender’ entity with randomly initialised, learnable features to serve as an actor, which is able to act on all nodes in the network.
@@
For the node-specific actions of Reduce Vulnerability and Restore Node, these can be implemented using either the ‘global’ categorical action space or the ‘entity-level’ categorical action space. For a GlobalCategoricalActionSpace, the RogueNet architecture creates a dummy global entity with randomly initialised learnable features, with information from the nodes of the network shared with its embedding vector through the attention mechanism in the Transformer layers. This single, compressed embedding is then provided as the input for the action head. Combined with the SelectEntityActionSpace, this means the agent chooses an ‘action type’ and then a node on which to execute it. Under the entity-level CategoricalActionSpace, all embeddings of the nodes are passed into the action head simultaneously in a batch, with a separate action selected for every node independently, based on each node’s respective embedding. Conceptually, combining this with a SelectEntityActionSpace would be like choosing a node to execute the action that it has selected, rather than choosing the action centrally. In our case, we chose to use the global action space to avoid computing redundant actions. Incidentally, the underlying mechanism is very similar to the approach outlined in Janisch et al., 2023b , where environments are represented as a graph, and an extra ‘global node’ is instantiated to gather global information over a series of GNN message-passing steps. This global node is then used to choose the ‘action identifier’, or action type, before proceeding in an autoregressive fashion.
@@ 6 Experiments
In this section, we outline the methodology and present results from the experiments we conducted to compare a baseline MLP-based Proximal Policy Optimization (PPO) agent (Schulman et al.,, 2017) with an entity-based RogueNet (Winter et al., 2023a, ) policy, as well as experiments exploring the zero-shot generalisation abilities of entity-based agents in environments with network topologies not seen during training. We drew the baseline PPO agent implementation from Stable Baselines 3 (Raffin et al.,, 2021), which is a library that collects stable and performant implementations of popular deep RL algorithms, and is generally reported to have the most consistent performance among open-source PPO implementations (Huang et al.,, 2024; Huang et al., 2022a, ). We trained the RogueNet agents using its companion Entity Neural Network Trainer package (Winter et al., 2023c, ) which uses a simple implementation of PPO derived from the one provided in the CleanRL library (Huang et al., 2022b, ).
Figure 2: Train-time episodic rewards evaluated on the 10, 20, and 40 node networks respectively. The four agents compared in each evaluation are: baseline PPO agent on the static (sb3_static_[nodes]) and random (sb3_random_[nodes]) network environments, and the entity neural network agent on the same static (sb3_random_[nodes]) and random (‘Entity_random_[nodes]’) environments. Rewards are averaged as the mean over three different random seeds, and shaded error bands are constructed between the maximum and minimum of the three runs. These bands are scarcely visible as there was not a lot of deviation between the three runs.
@@ 6.1 Training
As suggested in Mern et al., (2020, 2019), one advantage of the entity-based approach and using a Transformer policy is ‘object exchangeability’. That is, the robustness of Transformers and GNNs to different permutations of objects in the observation vector or action space without the effectiveness of the policy being impacted. To demonstrate this property, we consider two different training settings, which we refer to as Static and Random. In the Static setting, a single network is randomly initialised at the start of training, and remains the same for every episode; the only thing that varies is the identity of the entry node. In the Random setting, an entirely new network is initialised for each episode. The aim of this setup is to evaluate policy robustness across inhomogeneous networks in the training phase, and is similar to the method used by Hong et al., (2022) in ‘Multitask Inhomogeneous Reinforcement Learning’ for robotic control.
@@
To produce the training graphs, we used existing PPO hyperparameter configurations specified in the libraries for both agents. It should be noted that the learning rate used for the optimiser for the entity-based policy was 0.005, compared to 0.0003 in Stable Baselines 3. Different choices of learning rate are likely to affect the gap between asymptotic reward of the two varieties of policy in the Static training regime. The full configuration files for the training hyperparameters used can be found in the accompanying repository. A more rigorous analysis would involve hyperparameter sweeps for both varieties of agent, but we believe the fundamental conclusions around the performance difference between Random and Static modes, as well as the generalisability of entity-based agents, are robust.
@@
In total, we trained policies from 3 random seeds for each combination of policy type and environment setting. Considering all the combinations of agents, network sizes and ‘Random’ or ‘Static’ environments, we have trained 36 policies. Each policy was initialised and trained using a different random seed.
@@
  • sb3_static_$n$_node_network ×$\times 3$: Three different MLP policies trained using the Stable Baselines 3 PPO implementation, on static $n$-node networks randomly generated at the start of each training run, for $n\in(10,20,40)$ (the Static environment).
@@
  • sb3_random_$n$_node_network ×$\times 3$: Three different MLP policies trained using the Stable Baselines 3 PPO implementation, on $n$-node networks randomly generated with each episode reset for $n\in(10,20,40)$ (the Random environment).
  • EntityYT_static_$n$_node_network ×$\times 3$: Three different Entity Neural Network RogueNet policies, trained using the entity-based version of Yawning Titan, on static $n$-node networks randomly generated at the start of each training run, for $n\in(10,20,40)$ (the Static environment).
  • EntityYT_random_$n$_node_network ×$\times 3$: Three different Entity Neural Network RogueNet policies, trained using the entity-based version of Yawning Titan, on $n$-node networks randomly generated with each episode reset for $n\in(10,20,40)$ (the Random environment).
@@
Figure 2(a) shows the reward trajectories of all four types of policies trained on 10-node networks. Both entity-based agents (‘Entity_random_10’ and ‘Entity_static_10’) show similar reward trajectories that converge rapidly and plateau close to the maximum possible reward of 100. The Stable Baselines 3 PPO agents trained on the single static network (‘sb3_static_10’) converged quickly as well, to a slightly lower asymptotic reward. In line with expectations, the baseline PPO agents trained on a series of random networks that vary across episodes (‘sb3_random_10’) struggled to learn a good policy and displayed high variance in their performance. This supports the notion that a Transformer entity-based policy is more robust to exchangeable representations, and performs no worse than when trained on a static environment, whereas a fixed MLP-based policy struggles to learn an effective policy when the object that each element of the agent’s observation vector represents changes with each episode (Mern et al.,, 2020, 2019; Zhao et al.,, 2022).
@@
Figures 2(b) and (c) show similar results on 20-node and 40-node networks, respectively. Here, the difference between the asymptotic episodic reward of the policies trained in the Static regime grows smaller with a larger number of nodes. This may be, in part, due to the nature of the reward function defined as the proportion of non-compromised nodes at time $t$. Since the skill and behaviour of the red agent are unchanged, this naturally means that a lower proportion of nodes will be compromised on a given time step, and thus rewards for the blue agent will be higher in larger environments. Nevertheless, the Stable Baselines 3 agent trained on randomly generated networks still does comparatively poorly, and struggles to learn a policy that is invariant to different network topologies.